OpenStudy (anonymous):

3 dozen donuts to EACH of 15 local merchants. What is the total number of donuts sent?

OpenStudy (anonymous):

3*12*15=540

OpenStudy (anonymous):

thanks

OpenStudy (albert0898):

1 Dozen = 12 3 Dozen = 12 * 3 = 36 36 Donuts * 15 Merchants = 540 Donuts